The sum of the ages of a father and his son is 100 years now. 5 years ago their ages were in a the ratio of 2:1.. The ratio of the ages of father and son after 10 years will be?
Answer by jorel1380(3719)

Let s be the son. Then the father would be 100-s. And 2(s-5)=(100-s-5). So:
2s-10=100-s-5
3s=105
s=35
100-s=65
The son is 35. The father is 65. In 10 years their ratios will be 75:45, or 25:9. ☺☺☺☺